OpenStudy (anonymous):

Write the binary number \[ 11,101\overline{110}\] as decimal fraction and explain it with help of geometric series.

ganeshie8 (ganeshie8):

your comma means a binary point right ? (like decimal point)

OpenStudy (anonymous):

yes it is

OpenStudy (turingtest):

so they want, like, \[2^1+2^0+2^{-1}+0+2^{-3}+\sum k^{something}2^{something}\]? I don't think i get the question...

OpenStudy (turingtest):

maybe take the difference of two infinite sums to get those terms...

OpenStudy (anonymous):

@TuringTest yes that seems to be the gist. The given number can be written as \[2^1+2^0+2^{-1}+2^{-3}+\sum_{k=4}^\infty2^{-k}\] which can be summed nicely.

OpenStudy (anonymous):

Oops, the series isn't exactly right, but you get the idea.

OpenStudy (anonymous):

hm i have one smilar question with an answer it looks like \[0,\overline{1101}=\sum_{k=0}^{\infty} \frac{1}{2^{4k+1}}+\frac{1}{2^{4k+2}}+\frac{1}{2^{4k+4}}=...\]

OpenStudy (turingtest):

yeah we gotta subtract of another series to get those 0's in there i think

OpenStudy (turingtest):

\[\sum_{k=0}^\infty2^{2-k}-\sum_{k=0}^\infty2^{2-3k}\]? something like that?

OpenStudy (turingtest):

should be 4k\[\sum_{k=0}^\infty2^{2-k}-\sum_{k=0}^\infty2^{2-4k}\]not sure if that "explains" and I don't know how to write it as a "decimal fraction"

ganeshie8 (ganeshie8):

\[\large 11,101\overline{110} = 3.625+\sum_{k=0}^\infty2^{-(3k+4)} + 2^{-(3k+5)}\]

ganeshie8 (ganeshie8):

i think by decimal fraction they mean that sum for fractional part of binary number : 1/2 + 1/2^2 + ...

OpenStudy (anonymous):

yes ganeshie exactly

ganeshie8 (ganeshie8):

note that you can evaluate that infinite sum as |r| = 1/8 < 1